# The Name of the Title Is Hope

### v2

@@ 6.4. Verification on Edge Device
Although we theoretically calculate the peak memory footprint for model parameters, gradients, and intermediate feature maps, some real-applied consumption, such as extra consumption in loading datasets, memory fragmentation, and some additional intermediate variables, also need to be considered. Hence, we deploy LightCL  on an edge device, Jetson Nano, to verify the efficiency in the real environment. We compare LightCL  with SparCLDER++. As shown in Figure  5, although the peak memory of them is all greater than the theoretical value shown in Table 1, the peak memory of LightCL is still greatly smaller compared to SparCL. Note that the consumption of the memory footprint is not always equal to the peak memory footprint. We also need to observe the reserved memory, red line in Figure  5. Since the peak memory of LightCL  is smaller than SparCL, the reserved memory is also smaller than SparCL. From the x-axis that represents the change in time, the training process of LightCL  is faster than SparCL since the memory transfer consumption has greatly hidden the training speed in SparCL. All of these findings suggest the effectiveness of LightCL in edge devices.
Figure 5. We deploy LightCL (top figure) and SparCLDER++ (bottom figure) on Jetson Nano and evaluate the dynamic change of memory footprint during CL. Only $5$ iterations are selected.
@@ 7. Conclusion
This paper proposes an efficient CL  algorithm LightCL  to overcome the catastrophic forgetting problem in edge scenarios. It is the first study to explore the distribution of generalizability in neural networks by evaluating two designed metrics of learning plasticity and memory stability. This evaluation demonstrates that the generalizability of different layers in a neural network exhibits significant variation. Thus, we proposed LightCL  with Maintaining Generalizability and Memorize Feature Patterns. Experiments demonstrate that LightCL  is outstanding beyond other state-of-the-art   CL  methods with fewer resources, especially a lower memory footprint. Also, our method performs more favorably on edge devices and shows significant potential for efficient CL.

### v3

@@ 6.4. Verification on Edge Device
We deploy LightCL  on an edge device, Jetson Nano, to verify the efficiency in a realistic environment. Although we theoretically calculate the peak memory footprint for model parameters, gradients, and intermediate feature maps, some real-applied consumption, such as extra consumption in loading datasets, memory fragmentation, and some additional intermediate variables, also need to be considered. Hence, we conduct an experiment to compare LightCL  with SparCLDER++ on an edge device. As shown in Figure  5, although the peak memory of them is all greater than the theoretical value shown in Table 1, the peak memory of LightCL is still greatly smaller compared to SparCL. Note that the consumption of the memory footprint is not always equal to the peak memory footprint. We also need to observe the reserved memory, red line in Figure  5. Since the peak memory of LightCL  is smaller than SparCL, the reserved memory is also smaller than SparCL. From the x-axis that represents the change in time, the training process of LightCL  is faster than SparCL because the memory transfer consumption has greatly hindered the training speed in SparCL. These findings collectively demonstrate the effectiveness of LightCL on edge devices.
Figure 5. We deploy LightCL (top figure) and SparCLDER++ (bottom figure) on Jetson Nano and evaluate the dynamic change of memory footprint during CL. Only $5$ iterations are selected to illustrate this experiment.
@@ 7. Conclusion
This paper proposes a compact CL  algorithm LightCL  to overcome the catastrophic forgetting problem in edge scenarios. It is the first study to explore the distribution of generalizability in neural networks by evaluating two designed metrics of learning plasticity and memory stability. This evaluation demonstrates that the generalizability of different layers in a neural network exhibits a significant variation. Thus, we propose LightCL  with Maintaining Generalizability and Memorize Feature Patterns. Experiments demonstrate that LightCL  outpeforms other state-of-the-art   CL  methods with fewer resources, especially a lower memory footprint. Also, our method performs more favorably on edge devices and shows significant potential for efficient CL.
8. Acknowledgments
This work was supported by Educational Teaching Reform Practice Program No.A-XSY2308, University-Industry Collaborative Education Program No.231103428160213, Proof of Concept Foundation of Xidian University Hangzhou Institute of Technology under Grant No.GNYZ2023GY0301, and the Fundamental Research Funds for the Central Universities No.QTZX24064. Thanks to the help provided by the National Experimental Teaching Demonstration Center for Computer Network and Information Security affiliated with Xidian University. Also, we are grateful to Qichang Zhang for checking typos and grammar errors.